A drain has a rectangular profile. W is the width of the drain in meters, and


D is the depth of the water in meters.


In the case that the depth is much less than the width the amount of water that flows per second through a rectangular drain is given by the following formula:

Q =

A

n

D

2

3 S

1

2 (1)


In which:


- Q is the volume of water that


flows per second through the


drain (in m3/s);


- A = W×D, the cross-sectional


area of the drain up to the level


of the water (in m2)


- n is a parameter describing the


resistance to the flow of water;


- S is the gradient of the river


(in m/m).


Use formula 1 to derive the units of n.


1
Expert's answer
2021-08-23T09:10:43-0400

"Q = \frac{A}{n} D^{2/3} S^{1/2}"

Write equality for dimentions:

"[\frac{m^3}{s}] = \frac{[m^2]}{[n]} [m]^{2/3} [\frac{m}{m}]^{1/2}"


"[\frac{m^3}{s}] = \frac{[m^{8/3}]}{[n]}"


So units of n are "[n] = [\frac{s}{m^{1/3}}]"